OpenStudy (anonymous):

Solve by Factoring (Solve for SOLUTIONS): 2x^2+10x+8=0

OpenStudy (striker1298):

simplify the equation by dividing eveything by two

OpenStudy (striker1298):

that equals: \[x^2+5+4=0\]

OpenStudy (anonymous):

2x^2+10x+8=0 x^2+5x+4=0 (x +1) ( x+ 4)

OpenStudy (striker1298):

basically wat Gerry wrote

OpenStudy (anonymous):

Thank You all!

OpenStudy (anonymous):

you are very welcome

OpenStudy (striker1298):

no problem gerry deserves the medal more than me